### v1

# Parameter-related strong convergence rate and polynomial stability of a Euler’s type method for time-changed stochastic differential equations

Abstract
A Euler’s type method with the equidistant step size is proposed for a class of time-changed stochastic differential equations driven by the multiplicative noise and the strong convergence rate that is related to the parameter of the time changing process is obtained. Such a observation of the convergence rate is significantly different from those existing results that employ methods with the random step size. The polynomial stability in the mean square sense of the numerical method is also studied, which is in line with the asymptotic behavior of the underlying equation. Numerical simulations are provided to demonstrate the theoretical results.
keywords:
time-changed stochastic differential equations, equidistant-step Euler-Maruyama method, strong convergence, stability
1 Introduction
Stochastic differential equations (SDEs) driven by time-changed Brownian motion have emerged as a powerful mathematical framework for modeling anomalous diffusion phenomena across diverse scientific disciplines. These models have found widespread applications in physics [1], mathematical finance [2], and biological systems [3], where traditional diffusion models fail to capture the complex temporal patterns observed in experimental data. The key feature of time-changed SDEs is the incorporation of subordination processes, wherein a random time change governed by an inverse subordinator introduces memory effects and trapping mechanisms into the system dynamics [4]. This non-Markovian structure allows for more realistic modeling of systems exhibiting subdiffusive behavior, characterized by slower-than-normal diffusion rates. For a standard Itô SDE
$\mathrm{d}X(t)=f(X(t))\mathrm{d}t+g(X(t))\mathrm{d}B(t),$
the probability density function $p(x,t)$ of the solution $X(t)$ satisfies the Fokker-Planck equation(FPE)
$\frac{\partial p(x,t)}{\partial t}=\mathcal{L}_{\text{FP}}p(x,t),$
where $\mathcal{L}_{\text{FP}}=-\frac{\partial}{\partial x}\left(f(x)\cdot\right)+\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}\left(g(x)^{2}\cdot\right)$. In contrast, when considering subdiffusion phenomena via time-changed SDEs, as rigorously established by [5, 4], the probability density function of a time-changed process $X(E(t))$ (to be formally de?ned subsequently) is governed by a fractional Fokker-Planck equation(FFPE)
$D_{t}^{\alpha}p(x,t)=\mathcal{L}_{\text{FP}}p(x,t),$
where $D_{t}^{\alpha}$ represents the Caputo fractional derivative of order $\alpha$ with respect to $t$. However, obtaining analytical solutions to FFPE presents significant mathematical challenges. In this way, the solution of FFPE can be characterized through the probability density function of solutions to the corresponding time-changed SDEs.
Given the analytical complexity of time-changed SDEs, numerical methods play a crucial role in their practical implementation. The first systematic investigation of numerical methods for time-changed SDEs was conducted by Jum and Kobayashi [6], who considered the following time-changed SDEs
$\mathrm{d}Y(t)=f(E(t),Y(t))\mathrm{d}E(t)+g(E(t),Y(t))\mathrm{d}B(E(t)),\quad Y(0)=X_{0},$ (1)
where $E(t)$ represents the inverse subordinator and $B(E(t))$ denotes the time-changed Brownian motion. By establishing a duality principle that connects this time-changed SDEs to the classical Itô SDEs
$\mathrm{d}X(t)=f(t,X(t))\mathrm{d}t+g(t,X(t))\mathrm{d}B(t),\quad X(0)=X_{0},$ (2)
they showed that, under global Lipschitz conditions, if $X(t)$ solves eq. 2, then $Y(t)=X(E(t))$ solves eq. 1, and conversely, if $Y(t)$ solves eq. 1, then $X(t)=Y(D(t))$ solves eq. 2. Through this duality relationship, they developed an Euler-Maruyama (EM) method achieving a strong convergence rate of $1/2$.
Following this pioneering work, Jin and Kobayashi [7] investigated a class of time-changed SDEs where the inverse subordinator $E(t)$ in the drift and diffusion coefficients of eq. 1 is replaced with ordinary time $t$. In this modified framework, the duality principle no longer applies, necessitating the development of alternative analytical techniques. Using a Gronwall-like inequality specific to time-changed integrals, they demonstrated that the EM method maintains a strong convergence rate of $1/2$ in this setting as well.
Subsequently, the numerical methods of time-changed SDEs has been extensively studied under different discretization strategies. These include backward EM schemes [8], truncated EM methods [9, 10, 11], and higher-order methods such as the Milstein scheme [12, 13]. A notable commonality among these approaches is their reliance on random discretization of the inverse subordinator $E(t)$, which effectively transforms the stochastic increments of the time-change process into deterministic increments within the numerical scheme. As a consequence, these methods predominantly achieve convergence rates similar to those established for classical SDEs (typically of rate $1/2$ or $1$), despite the additional complexity introduced by the time-change mechanism.
A natural question arises: can we develop numerical schemes for time-changed SDEs whose convergence properties explicitly reflect the unique characteristics of the time-change process? This question motivates the central innovation of our present work: an equidistant-step EM method for time-changed SDEs. Unlike previous approaches, our method preserves the intrinsic stochasticity of the subordinator process increments within the numerical scheme, leading to convergence behavior that directly depends on the stability index $\alpha$ of the subordinator.
Stability of stochastic systems is also an active research area[14]. Most existing studies of stability for time-changed SDEs focused on the following form
$\mathrm{d}X(t)=h(t,E(t),X(t))\mathrm{d}t+f(t,E(t),X(t))\mathrm{d}E(t)+g(t,E(t),X(t))\mathrm{d}B(E(t)).$ (3)
Wu [15] established stochastic stability, stochastic asymptotic stability, and global stochastic asymptotic stability of time-changed SDEs. Zhu et al. [16] demonstrated almost sure exponential stability under appropriate coefficient conditions, while subsequent work by Zhu et al. [17] established exponential stability, revealing the term $h(t,E(t),X(t))\mathrm{d}t$ in eq. 3 determining these stability properties. Furthermore, the stability theory for time-changed stochastic systems has been extended to more general noise structures by Nane and Ni [18, 19], who established comprehensive stability results for SDEs driven by time-changed Lévy noises. This direction was further developed by Yin et al. [20], who investigated the stability properties of such systems under impulsive effects.
However, as noted in Remark 2.13 of [19], the inverse subordinator $E(t)$, which will be formally introduced in Section 2, exhibits asymptotically slower growth compared to ordinary time $t$. Specifically, it satisfies the asymptotic relation $\lim_{t\to\infty}\frac{E(t)}{t}=0$ almost surely. Therefore, the term $h(t,E(t),X(t))\mathrm{d}t$ tends to dominate the system dynamics, potentially masking the distinctive effects introduced by the time-change mechanism. To isolate and analyze more precisely the impact of time-change on stability properties, Deng and Liu [8] and Liu [21] studied the following time-changed SDEs
$\mathrm{d}X(t)=f(t,E(t),X(t))\mathrm{d}E(t)+g(t,E(t),X(t))\mathrm{d}B(E(t)),$ (4)
in which the system dynamics are governed exclusively by the time-changed components.
Within this more focused framework, Deng and Liu [8] established a profound connection between the polynomial stability index of time-changed SDEs and the stability parameter of the subordinator. Their analysis extended to numerical methods as well, demonstrating polynomial stability for the backward EM scheme. Complementing these findings, Liu [21] investigated polynomial stability under polynomial growth conditions, confirming that this stability property is also fundamentally characterized by the stability parameter of the subordinator.
For analytical simplicity in this paper, we concentrate on a simplified form of (4), specifically
$\mathrm{d}X(t)=f(X(t))\mathrm{d}E(t)+g(X(t))\mathrm{d}B(E(t)),$ (5)
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ are coefficient functions satisfying global Lipschitz and linear growth conditions, $E(t)$ denotes the inverse subordinator, and $B(E(t))$ represents the $m$-dimensional time-changed standard Brownian motion.
Our contributions are threefold
  1. 1.
    We develop an equidistant-step EM method for time-changed SDEs and establish its strong convergence properties under global Lipschitz conditions.
  2. 2.
    We rigorously prove that the strong convergence rate of our method is $\mathcal{O}(\Delta t^{(1+\alpha)/4})$, explicitly characterizing how the stability index $\alpha$ of the subordinator process determines the convergence behavior. This result reveals a fundamental relationship between the statistical properties of the time-change mechanism and the accuracy of the numerical method.
  3. 3.
    We conduct a comprehensive stability analysis of the numerical scheme, deriving sufficient conditions for asymptotic mean-square polynomial stability with decay rate proportional to $t^{-\alpha}$, thus establishing a direct link between the memory effects induced by the time-change process and the long-term behavior of the numerical solution.
The remainder of this paper is organized as follows Section 2 introduces the mathematical preliminaries, including key properties of subordinators and their inverses, and establishes important lemmas used in our subsequent analysis. Section 3 presents the fundamental assumptions, develops the equidistant-step EM scheme, and proves our main convergence theorem. Section 4 investigates the stability properties of both the exact and numerical solutions. Section 5 presents numerical experiments that validate our theoretical findings and illustrate the effectiveness of the proposed method.
2 Preliminaries
Throughout this paper, we work within a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. We begin by introducing the fundamental stochastic processes that underlie our analysis of time-changed SDEs.
Let $D=(D(t))_{t\geq 0}$ denote a non-decreasing Lévy process with $D(0)=0$, characterized by cádĺág (right-continuous with left limits) sample paths. Such a process $D$ is called a subordinator and is uniquely determined by its Laplace exponent $\psi:[0,\infty)\to[0,\infty)$ through the relation
$\mathbb{E}[e^{-sD(t)}]=e^{-t\psi(s)},\quad s\geq 0,\quad t\geq 0.$
The Laplace exponent $\psi$, with zero killing rate, is represented via the Lévy-Khintchine formula as
$\psi(s)=\int_{0}^{\infty}(1-e^{-sy})\,\nu(\mathrm{d}y),\quad s>0,$
where $\nu$ is the Lévy measure satisfying the integrability condition $\int_{0}^{\infty}(y\wedge 1)\,\nu(\mathrm{d}y)<\infty$. Throughout this paper, we assume that $\nu(0,\infty)=\infty$, which ensures that $D$ has strictly increasing paths with infinitely many jumps. The inverse subordinator $E=(E(t))_{t\geq 0}$ associated with $D$ is defined by
$E(t):=\inf\{u>0:D(u)>t\},\quad t\geq 0.$
This process $E$ is continuous, non-decreasing, and initialized at $E(0)=0$. The fundamental relationship between $D$ and $E$ is characterized by the equivalence
$\{E(t)>x\}=\{D(x)<t\},\quad\forall t,x\geq 0.$
In this work, we focus specifically on the case where $D$ is an $\alpha$-stable subordinator with stability index $\alpha\in(0,1)$, for which the Laplace exponent takes the simplified form $\psi(s)=s^{\alpha}$. Consequently, $E$ is referred to as the inverse $\alpha$-stable subordinator. The assumption $\nu(0,\infty)=\infty$ ensures that $E$ has continuous sample paths. A critical feature of this process is that jumps in $D$ correspond to intervals of constancy in $E$, during which any time-changed process $X\circ E=(X(E(t)))_{t\geq 0}$ remains constant.
Let $B=(B(t))_{t\geq 0}$ be a standard $m$-dimensional Brownian motion defined on $(\Omega,\mathcal{F},\mathbb{P})$ and independent of the subordinator $D$. We denote by $\mathbb{P}_{B}$ and $\mathbb{P}_{D}$ the probability measures associated with $B$ and $D$, respectively, and by $\mathbb{E}_{B}$ and $\mathbb{E}_{D}$ the corresponding expectation operators. The total expectation $\mathbb{E}$ under the probability measure $\mathbb{P}$ can then be expressed as
$\mathbb{E}[\cdot]=\mathbb{E}_{D}[\mathbb{E}_{B}[\cdot]]=\mathbb{E}_{B}[\mathbb{E}_{D}[\cdot]].$
The time-changed process $B\circ E=(B(E(t)))_{t\geq 0}$ represents a Brownian motion evaluated along the random time scale defined by $E$. This construction effectively models particles that remain immobile during the constant intervals of $E$, capturing the characteristic trapping phenomena observed in subdiffusive systems. It is important to note that while $B\circ D$ inherits the Brownian motion property from $B$, the time-changed process $B\circ E$ is non-Markovian, reflecting the memory effects introduced by the inverse subordinator. The quadratic variations of these processes satisfy the following relationships
$B\circ E,B\circ E]_{t}=E(t),\quad[E,E]_{t}=[B\circ E,E]_{t}=0,\quad t\geq 0.$
For a comprehensive treatment of stochastic calculus for time-changed semimartingales, we refer the reader to [4]. Figure 1 provides a visual representation of the processes $D$, $E$, and $B\circ E$.
Figure 1: The left figure is sample paths of an 0.8-stable subordinator $D$ (blue) and its inverse subordinator $E$ (red). The right figure is sample paths of an inverse 0.8-stable subordinator $E$ (blue) and the corresponding time-changed Brownian motion $B\circ E$ (red).
To numerically simulate the inverse subordinator $E$, we employ the method developed by Magdziarz [22, 23]. Given a time horizon $T>0$ and an equidistant step size $\delta>0$, we first generate a sample path of $D$ over $[0,T]$ by setting $D(0)=0$ and recursively computing
$D(i\delta)=D((i-1)\delta)+Z_{i},\quad i=1,2,\ldots,$
where $(Z_{i})_{i\in\mathbb{N}}$ constitutes an independent and identically distributed sequence with $Z_{i}\sim\Delta D_{\delta}$. These increments are generated using the formula
$Z_{i}=\frac{\sin(\alpha(V+\pi/2))}{(\cos(V))^{1/\alpha}}\left(\frac{\cos(V-\alpha(V+\pi/2))}{W}\right)^{(1-\alpha)/\alpha},$
where $V\sim\text{Uniform}(-\pi/2,\pi/2)$ and $W\sim\text{Exponential}(1)$ are independent random variables. The approximation $\tilde{E}=(\tilde{E}(t))_{t\geq 0}$ of the inverse subordinator is then defined by
$\tilde{E}(t):=\left(\min\{n\in\mathbb{N}:D(n\delta)>t\}-1\right)\delta,\quad t\geq 0.$
This approximation $\tilde{E}$ is a non-decreasing step function with jump size $\delta$. As demonstrated in [22, 23], it provides a close approximation to $E$ with the error bounds
$E(t)-\delta\leq\tilde{E}(t)\leq E(t),\quad\text{for all }t\in[0,T],$
holding almost surely. For the time-changed SDEs eq. 5, we consider the equidistant-step EM method. Given a uniform partition $\{t_{n}\}_{n=0}^{N}$ of $[0,T]$ with step size $\Delta t=t_{n+1}-t_{n}$, the EM scheme is defined recursively by
$X_{n+1}=X_{n}+f(X_{n})\Delta E_{n}+g(X_{n})\Delta B_{n},\quad n=0,1,\ldots,N-1,$ (6)
where $\Delta E_{n}=E(t_{n+1})-E(t_{n})$ and $\Delta B_{n}=B(E(t_{n+1}))-B(E(t_{n}))$ represent the increments of the inverse subordinator and time-changed Brownian motion, respectively. To facilitate theoretical analysis, we introduce a piecewise constant interpolation of the numerical solution, defined by
$\tilde{X}_{t}=X_{n}\quad\text{for }t\in[t_{n},t_{n+1}),\quad n=0,1,\ldots,N-1.$
For any $s\in[0,T]$, let $n_{s}$ denote the index such that $s\in[t_{n_{s}},t_{n_{s}+1})$. The discrete scheme eq. 6 can then be expressed in the equivalent integral form
$\tilde{X}_{t_{n_{s}}}=X_{0}+\int_{0}^{t_{n_{s}}}f(\tilde{X}_{r})\,\mathrm{d}E(r)+\int_{0}^{t_{n_{s}}}g(\tilde{X}_{r})\,\mathrm{d}B(E(r)),$ (7)
which provides a natural framework for analyzing the convergence properties and error behavior of the EM method.
The equidistant-step EM method presented in eq. 6 has been previously utilized by Carnaffan and Kawai [24] in their investigation of anomalous diffusion processes, which provide stochastic representations for multidimensional FFPE. Their approach employed sophisticated techniques from Malliavin calculus to derive unbiased density formulas for the probability density functions of these anomalous diffusion processes. This methodology yielded remarkably accurate numerical approximations that successfully captured essential characteristics of the underlying processes, including the distinctive cusp formations observed in density profiles. Despite these promising applications, a rigorous mathematical analysis concerning the convergence properties and stability characteristics of this numerical scheme in the context of time-changed SDEs has not been thoroughly established in the literature. The present work endeavors to fill this significant gap in the theoretical understanding of numerical methods.
To establish our theoretical results, we require several concepts from the theory of regularly varying functions, which play a crucial role in characterizing the asymptotic behavior of processes involving the inverse subordinator. A function $f:(0,\infty)\to(0,\infty)$ is said to be regularly varying at infinity with index $\beta\in\mathbb{R}$ if for any $c>0$,
$\lim_{s\to\infty}\frac{f(cs)}{f(s)}=c^{\beta}.$
We denote by $\text{RV}_{\beta}$ the class of regularly varying functions at infinity with index $\beta$. A function is called slowly varying at infinity if it belongs to $\text{RV}_{0}$. Every regularly varying function $f$ with index $\beta$ admits a representation $f(s)=s^{\beta}\ell(s)$, where $\ell$ is slowly varying.
In the subsequent analysis, we employ several key results concerning regularly varying functions and their applications to the moments of the inverse subordinator. We now present two fundamental lemmas from [12, 7] that constitute the theoretical cornerstone of our convergence analysis.
Lemma 1.
The following properties hold for regularly varying functions.
  1. (i)
    Given $f\in\textup{RV}_{p}(\infty)$ (a regularly varying function at infinity with index $p$), $f(\infty)=\infty$ if $p>0$, and $f(\infty)=0$ if $p<0$.
  2. (ii)
    If $f_{i}\in\textup{RV}_{p_{i}}(\infty)$ for $i=1,2$ and $f_{2}(\infty)=\infty$, then the composition $f_{1}\circ f_{2}\in\textup{RV}_{p_{1}p_{2}}(\infty)$.
  3. (iii)
    If $f_{i}\in\textup{RV}_{p_{i}}(\infty)$ for $i=1,2$, then the product $f_{1}\cdot f_{2}\in\textup{RV}_{p_{1}+p_{2}}(\infty)$.
Lemma 2.
Let $E$ be the inverse of a subordinator $D$ whose Laplace exponent $\psi$ is regularly varying at $\infty$ with index $\alpha\in[0,1)$. If $\alpha=0$, assume further that $\psi(\infty)=\infty$ and $\psi^{\prime}$ is regularly varying at $\infty$ with index 1-1. Suppose $f:(0,\infty)\to(0,\infty)$ is a measurable function regularly varying at $\infty$ with index $p>0$ such that $\sup_{s\leq s_{0}}f(s)<\infty$ for any $0<s_{0}<\infty$. Fix $t>0$.
  1. (i)
    If $p<1/(1-\alpha)$, or equivalently, $\alpha>(p-1)/p$, then $\mathbb{E}[e^{f(E_{t})}]<\infty$.
  2. (ii)
    If $p>1/(1-\alpha)$, or equivalently, $\alpha<(p-1)/p$, then $\mathbb{E}[e^{f(E_{t})}]=\infty$.
Lemma 2 establishes important criteria for the finiteness of exponential moments of the inverse subordinator process $E$. This result plays a crucial role in our analysis by providing precise conditions under which certain expectations remain bounded, directly affecting the convergence properties of our numerical scheme. To illustrate the significance of this result, consider the specific case of the inverse $\alpha$-stable subordinator $E$ with Laplace exponent $\psi(s)=s^{\alpha}$, where $\alpha\in(0,1)$. For the function $f(E(t))=E^{2}(t)$, applying Lemma 2 reveals a critical threshold phenomenon: $\mathbb{E}\left[e^{f(E(t))}\right]=\infty$ when $\alpha\in(0,\frac{1}{2})$, whereas $\mathbb{E}\left[e^{f(E(t))}\right]<\infty$ when $\alpha\in(\frac{1}{2},1)$. This explicit characterization demonstrates how the stability index $\alpha$ fundamentally determines the existence of exponential moments, a property that will be essential in our subsequent convergence analysis.
The following lemma, which can also be found in [25], provides precise bounds on the expected values of multiple integrals involving the inverse subordinator. This result constitutes the cornerstone of our convergence rate analysis, establishing the direct relationship between the time step size and the expectation of multiple integrals driven by increments of $E(t)$.
Lemma 3.
Let $E$ be the inverse process of a subordinator $D$ with stability index $\alpha\in(0,1)$. For any fixed $\varepsilon>0$ and $a,b$ satisfying $\varepsilon<a<b$, and for any positive integer $n$, the following bound holds
$\mathbb{E}\left[\left(\int_{a}^{b}\mathrm{d}E(t)\right)^{n}\right]\leq C(b-a)^{1+(n-1)\alpha},$
where $C>0$ is a constant independent of bab-a.
Proof.
The proof follows from similar arguments as in [25]. For completeness, we provide a detailed proof in A. ∎
While Lemma 3 provides a general bound for arbitrary powers $n$, it is instructive to examine the special case $n=1$ more closely. Below, we derive an explicit formula for the expected increment of the inverse $\alpha$-stable subordinator over an arbitrary interval $[a,b]$ with $0<a<b$, which complements and illuminates the general result.
Utilizing the well-established moment formula $\mathbb{E}\left[E^{n}(t)\right]=\frac{n!t^{n\alpha}}{\Gamma(1+n\alpha)}$ for $n\geq 1$, we derive the first moment of the increment as follows
$\mathbb{E}\left[E(b)-E(a)\right]$ $=\mathbb{E}\left[E(b)\right]-\mathbb{E}\left[E(a)\right]=\frac{b^{\alpha}-a^{\alpha}}{\Gamma(1+\alpha)}.$
By applying the mean value theorem for differentiable functions, we can identify a point $\xi\in(a,b)$ such that
$b^{\alpha}-a^{\alpha}=\alpha\xi^{\alpha-1}(b-a).$
Consequently, we obtain the elegant representation
$\mathbb{E}\left[E(b)-E(a)\right]=\frac{\alpha\xi^{\alpha-1}}{\Gamma(1+\alpha)}(b-a),\quad\text{where }\xi\in(a,b).$
For any fixed $\varepsilon>0$, when we consider $a\geq\varepsilon$ (ensuring we avoid the singularity at the origin), we have $\xi^{\alpha-1}\leq\varepsilon^{\alpha-1}$ for all $\xi\geq\varepsilon$ (since $\alpha<1$). This yields the important upper bound
$\mathbb{E}\left[E(b)-E(a)\right]\leq C(b-a),$
where $C$ is a positive constant depending only on $\alpha$ and $\varepsilon$.
The moment bounds established in this section form the cornerstone of our subsequent analysis of numerical methods for time-changed SDEs. They provide rigorous control over the stochastic increments driven by the inverse subordinator $E(t)$ and directly influence the achievable convergence rates of our numerical scheme.
3 The strong convergence of EM method
Assumption 4.
There exists a constant $K$ such that for all d $x,y\in\mathbb{R}^{d}$,
$|f(x)-f(y)|+|g(x)-g(y)|\leq K|x-y|.$
Assumption 5.
There exists a constant $K$ such that for all $x\in\mathbb{R}^{d}$,
$|f(x)|+|g(x)|\leq K(1+|x|).$
These two assumptions ensure the existence and uniqueness of the solution to the SDEs eq. 5. The following proposition, adapted from Lemma 8 in [7] , establishes the moment boundedness of the true solution to the time-changed SDEs eq. 5.
Proposition 6.
Let $X$ be the solution of SDEs eq. 5, where $f$ satisfies Assumption 4 and Assumption 5. Then for any $p\geq 1$,we have $\mathbb{E}[Y^{(p)}(T)]<\infty$, where $Y^{(p)}(t)=1+\sup_{0\leq s\leq t}|X(s)|^{p}$.
We now turn our attention to the main focus of this paper, namely the strong convergence properties of the equidistant-step EM method. Before proceeding, we recall the notion of strong convergence for numerical schemes. A numerical method $X_{\Delta t}$ of a stochastic process $X$ is said to converge strongly with order $\gamma>0$ if there exists a constant $C$, independent of the step size $\Delta t$, such that
$\mathbb{E}\left[\sup_{t\in[0,T]}|X(t)-X_{\Delta t}(t)|\right]\leq C(\Delta t)^{\gamma},$
where $T$ is the terminal time.
Theorem 7.
Let $\tilde{X}$ be the approximation process defined in eq. 7 for the true solution X of SDEs eq. 5. Under the Assumptions 4 and 5, and assuming that $\alpha\in(1/2,1)$ is the stability index of the subordinator process $D$, there exists a constant $C$ independent of $\Delta t$ such that
$\mathbb{E}\left[\sup\limits_{t\in[0,T]}|X({t})-\tilde{X}_{t}|\right]\leq C\Delta t^{\frac{1+\alpha}{4}}.$
Proof.
To measure the deviation between the true solution of the SDEs and the numerical method, we introduce the error term
$Z(t):=\sup_{s\in[0,t]}|X(s)-\tilde{X}_{n_{s}}|,$
where $X(s)$ represents the true solution at time $s$ and $\tilde{X}_{n_{s}}$ is the numerical method at time $t_{n_{s}}$.
By comparing eq. 5 andeq. 7, we obtain
$X(s)-\tilde{X}_{n_{s}}$ $=\int_{0}^{t_{n_{s}}}f(X(r))-f(\tilde{X}_{r})\,\mathrm{d}E(r)+\int_{0}^{t_{n_{s}}}g(X(r))-g(\tilde{X}_{r})\,\mathrm{d}B(E(r))$
$\quad+\int_{t_{n_{s}}}^{s}f(X(r))\,\mathrm{d}E(r)+\int_{t_{n_{s}}}^{s}g(X(r))\,\mathrm{d}B(E(r)),$
where $t_{n_{s}}$ denotes the discrete time point such that $s\in[t_{n_{s}},t_{n_{s}+1})$. From this decomposition, we derive the following bound for the error
$Z(t)\leq I_{1}+I_{2}+I_{3}+I_{4},\quad Z^{2}_{t}\leq 4(I^{2}_{1}+I^{2}_{2}+I^{2}_{3}+I^{2}_{4}),$
where the terms are defined as
$I_{1}$ $=\sup_{s\in[0,t]}\left|\int_{0}^{t_{n_{s}}}f(X(r))-f(\tilde{X}_{r})\,\mathrm{d}E(r)\right|,$
$I_{2}$ $=\sup_{s\in[0,t]}\left|\int_{0}^{t_{n_{s}}}g(X(r))-g(\tilde{X}_{r})\,\mathrm{d}B(E(r))\right|,$
$I_{3}$ $=\sup_{s\in[0,t]}\left|\int_{t_{n_{s}}}^{s}f(X(r))\,\mathrm{d}E(r)\right|,$
$I_{4}$ $=\sup_{s\in[0,t]}\left|\int_{t_{n_{s}}}^{s}g(X(r))\,\mathrm{d}B(E(r))\right|.$
For $I_{1}$, by Assumption 4, we have
$I_{1}$ $\leq\sup_{s\in[0,t]}\int_{0}^{t_{n_{s}}}\big|f(X(r))-f(\tilde{X}_{r})\big|\,\mathrm{d}E(r)$
$\leq\sup_{s\in[0,t]}\int_{0}^{t_{n_{s}}}K|X(r)-\tilde{X}_{r}|\,\mathrm{d}E(r)$
$\leq K\int_{0}^{t_{n_{t}}}Z(r)\,\mathrm{d}E(r),$
Since $t_{n_{t}}\leq t\leq T$, using Cauchy-Schwarz inequality and taking the expectation $\mathbb{E}_{B}$ on both sides, we have
$\mathbb{E}_{B}\left[I_{1}^{2}\right]$ $\leq K^{2}E(T)\int_{0}^{t}\mathbb{E}_{B}[Z^{2}(r)]\,\mathrm{d}E(r).$ (8)
For $I_{2}$, by Burkholder-Davis-Gundy(BDG) inequality and Assumption 4, we have
$\mathbb{E}_{B}[I_{2}^{2}]$ $\leq\mathbb{E}_{B}\left[\int_{0}^{t_{n_{t}}}\big|g(X(r))-g(\tilde{X}_{r})\big|^{2}\,\mathrm{d}E(r)\right]$
$\leq\mathbb{E}_{B}\left[\int_{0}^{t_{n_{t}}}K^{2}|X(r)-\tilde{X}_{r}|^{2}\,\mathrm{d}E(r)\right]$
$\leq K^{2}\int_{0}^{t_{n_{t}}}\mathbb{E}_{B}[Z^{2}(r)]\,\mathrm{d}E(r).$ (9)
For $I_{3}$, by Assumption 5, we have
$I_{3}$ $\leq\sup_{s\in[0,t]}\int_{t_{n_{s}}}^{s}\left|f(X(r))\right|\,\mathrm{d}E(r)$
$\leq\sup_{s\in[0,t]}\int_{t_{n_{s}}}^{s}K\big(1+|X(r)|\big)\,\mathrm{d}E(r)$
$\leq\max_{k}\int_{t_{k}}^{t_{k+1}\wedge t}KY^{(1)}(r)\,\mathrm{d}E(r),$
where $k=0,1,\ldots,n_{t}$. Then, by Hölder’s inequality for the measure $\mathrm{d}E(r)$, we have
$I_{3}^{4}$ $\leq\sum_{k=0}^{n_{t}}\left(\int_{t_{k}}^{t_{k+1}\wedge t}KY^{(1)}(r)\,\mathrm{d}E(r)\right)^{4}$
$\leq\sum_{k=0}^{n_{t}}(\Delta E_{k})^{3}\int_{t_{k}}^{t_{k+1}\wedge t}K^{4}Y^{(4)}(r)\,\mathrm{d}E(r)$
Taking the expectation $\mathbb{E}_{B}$ on both sides, we have
$\mathbb{E}_{B}[I_{3}^{4}]$ $\leq K^{4}\sum_{k=0}^{n_{t}}(\Delta E_{k})^{3}\int_{t_{k}}^{t_{k+1}\wedge t}\mathbb{E}_{B}[Y^{(4)}(r)]\,\mathrm{d}E(r)$
$\leq K^{4}\mathbb{E}_{B}[Y^{(4)}(T)]\sum_{k=0}^{n_{t}}(\Delta E_{k})^{4}.$ (10)
Thus, we have
$\mathbb{E}_{B}[I_{3}^{2}]$ $\leq K^{2}\left[\mathbb{E}_{B}[Y^{(4)}(T)]\sum_{k=0}^{n_{t}}(\Delta E_{k})^{4}\right]^{\frac{1}{2}}.$
For $I_{4}$, by BDG inequality and Assumption 5, we have
$\mathbb{E}_{B}[I_{4}^{2}]$ $\leq\mathbb{E}_{B}\left[\int_{t_{n_{t}}}^{t}\left|g(E(r),X(r))\right|^{2}\,\mathrm{d}E(r)\right]$
$\leq\mathbb{E}_{B}\left[\int_{t_{n_{t}}}^{t}K^{2}\big(1+|X(r)|^{2}\big)\,\mathrm{d}E(r)\right]$
$\leq K^{2}\int_{t_{n_{t}}}^{t}\mathbb{E}_{B}[Y^{(2)}(r)]\,\mathrm{d}E(r)$
$\leq K^{2}\mathbb{E}_{B}[Y^{(2)}(T)](E(t)-E(t_{n_{t}})).$ (11)
Combining eqs. 8 to 3, we obtain
$\mathbb{E}_{B}[Z^{2}(t)]$ $\leq 4K^{2}(E(T)+1)\int_{0}^{t}\mathbb{E}_{B}[Z^{2}(r)]\,\mathrm{d}E(r)$
$\quad+4K^{2}\mathbb{E}_{B}[Y^{(2)}(T)]\left(E(t)-E(t_{n_{t}})\right)+4K^{2}\left[\mathbb{E}_{B}[Y^{(4)}(T)]\sum_{k=0}^{n_{t}}(\Delta E_{k})^{4}\right]^{\frac{1}{2}}.$
Applying Gronwall’s inequality, we get
$\mathbb{E}_{B}[Z^{2}(t)]$ $\leq\left(4K^{2}\mathbb{E}_{B}[Y^{(2)}(T)]\left(E(t)-E(t_{n_{t}})\right)+4K^{2}\left[\mathbb{E}_{B}[Y^{(4)}(T)]\sum_{k=0}^{n_{t}}(\Delta E_{k})^{4}\right]^{\frac{1}{2}}\right)$
$\quad\cdot e^{4K^{2}(E(T)+1)E(T)}.$
Taking the expectation $\mathbb{E}_{D}$ on both sides and using Cauchy-Schwarz inequality, we have
$\mathbb{E}[Z^{2}(t)]$ $\leq 4K^{2}\left(\mathbb{E}[Y^{(8)}(T)]\right)^{\frac{1}{4}}\left(\mathbb{E}\left[e^{16K^{2}(E(T)+1)E(T)}\right]\right)^{\frac{1}{4}}$
$\quad\cdot\left[\left(\mathbb{E}\left[E(t)-E(t_{n_{t}})\right]^{2}\right)^{\frac{1}{2}}+\left(\sum_{k=0}^{n_{t}}\mathbb{E}[(\Delta E_{k})^{4}]\right)^{\frac{1}{2}}\right]$
By Proposition 6, we have $\mathbb{E}[Y^{(8)}(T)]<\infty$. For the exponential term $\mathbb{E}\left[e^{16K^{2}(E(T)+1)E(T)}\right]$, since $\alpha>0.5$, Lemmas 1 and 2 guarantee this expectation is finite. For the term involving the increments of $E(t)$, applying Lemma 3 and noting that $t-t_{n_{t}}\leq\Delta t$, we obtain
$\left(\mathbb{E}\left[E(t)-E(t_{n_{t}})\right]^{2}\right)^{\frac{1}{2}}+\left(\sum_{k=0}^{n_{t}}\mathbb{E}[(\Delta E_{k})^{4}]\right)^{\frac{1}{2}}\leq C\left(\Delta t^{(1+\alpha)}+\Delta t^{3\alpha}\right)^{\frac{1}{2}}$
where $C$ is a positive constant independent of $\Delta t$.
Therefore, we derive
$\mathbb{E}[Z(t)]$ $\leq C\left(\Delta t^{1+\alpha}+\Delta t^{3\alpha}\right)^{\frac{1}{4}}$
$\leq C\Delta t^{\frac{1+\alpha}{4}},$
where the last inequality follows from the fact that $\alpha>0.5$. This completes the proof. ∎
Remark 1.
As the convergence rate $(1+\alpha)/4$ indicates, when $\alpha$ approaches $1$, the convergence rate tends to the standard rate of $1/2$, which coincides with the convergence rates established by [6, 8] and others for non-equidistant step size methods. Furthermore, if we consider subordinators with a Laplace exponent that includes a positive linear drift term, as studied in [8], it can be shown that there exists a finite constant $C$ such that $\Delta E_{k}<C\Delta t$ for all $k$. In this case, all $\Delta E_{k}$ terms in Theorem 7 can be bounded by $\Delta t$, leading to the classical convergence rate of $1/2$. We will verify this observation numerically in Section 5.
4 Stability analysis of the EM method
Having established the strong convergence properties of the EM method, we now proceed to investigate its stability characteristics. We begin by giving the definitions of asymptotic mean square polynomial stability for the true solution and the numerical solution.
Definition 1.
The true solution $X(t)$ of the time-changed SDEs eq. 5 is said to be asymptotically polynomially stable in the mean square sense if there exists a positive constant $p$ such that
$\limsup_{t\to\infty}\frac{\log\mathbb{E}\left[|X(t)|^{2}\right]}{\log t}\leq-p.$
Definition 2.
The EM method defined by eq. 6 is said to be asymptotically polynomially stable in the mean square sense if there exists a positive constant $p$ such that
$\limsup_{n\to\infty}\frac{\log\mathbb{E}\left[|X_{n}|^{2}\right]}{\log t_{n}}\leq-p.$
To facilitate our stability analysis, we introduce the following assumptions on the coefficients of the time-changed SDEs.
Assumption 8.
There exists a positive constant $K$ such that for all $x\in\mathbb{R}^{d}$,
$|f(x)|\leq K|x|.$
Assumption 9.
There exist positive constants $L$ such that for all $x\in\mathbb{R}^{d}$,
$2\langle x,f(x)\rangle+|g(x)|^{2}\leq-L|x|^{2}.$
The condition in Assumption 9 is a strong dissipativity condition, which ensures that the combined effect of the drift term and the squared diffusion coefficient is negative definite. This guarantees that the system has a strong stabilizing tendency toward the origin, with the drift term’s contractive effect dominating any potential destabilizing influence from the diffusion term. This type of condition is commonly used in the stability analysis of SDEs and their numerical methods.
We now present our main theorem regarding the asymptotic stability properties of both the true solution and the EM method for time-changed SDEs.
Theorem 10.
Let $E$ be the inverse process of the subordinator $D$ with stability index $\alpha\in(0,1)$. Under Assumption 9, $f$ the initial condition $X_{0}$ is a constant, then the true solution of the time-changed SDEs eq. 5 is asymptotically polynomially stable in the mean square sense, i.e.,
$\limsup_{t\to\infty}\frac{\log\mathbb{E}\left[|X(t)|^{2}\right]}{\log t}\leq-\alpha.$
To prove this theorem, we require the following lemma from [8], which characterizes the asymptotic behavior of the inverse subordinator.
Lemma 11.
If the Bernstein function $\phi\in\textup{RV}_{\alpha}$ with $\alpha\in(0,1)$, then for any $\lambda>0$,
$\lim_{t\to\infty}\frac{\log\mathbb{E}_{D}\left[e^{-\lambda E(t)}\right]}{\log t}=-\alpha.$
Proof of Theorem 10.
The proof follows from similar arguments as in [21]. For completeness, we provide a detailed proof in B. ∎
Now, we proceed to analyze the stability properties of the EM method.
Theorem 12.
Let $E$ be the inverse process of the subordinator $D$ with stability index $\alpha\in(0,1)$. Under Assumptions 8 and 9, consider a family of uniform grids with step size $\Delta t\to 0$ and $t_{n}=n\Delta t\to\infty$. If the initial condition $X_{0}$ is a constant, then the EM method exhibits asymptotic polynomial stability in the mean-square sense with precise decay rate $\alpha$, i.e.,
$\limsup_{n\to\infty}\frac{\log\mathbb{E}\left[|X_{n}|^{2}\right]}{\log t_{n}}\leq-\alpha.$
Proof.
We begin by taking the square of the Euclidean norm on both sides of the iteration formula in eq. 6, which yields
$|X_{k+1}|^{2}$ $=|X_{k}|^{2}+|f(X_{k})|^{2}|\Delta E_{k}|^{2}+|g(X_{k})|^{2}|\Delta B_{k}|^{2}+2\langle X_{k},f(X_{k})\rangle\Delta E_{k}$
$\quad+2\langle X_{k},g(X_{k})\Delta B_{k}\rangle+2\langle f(X_{k})\Delta E_{k},g(X_{k})\Delta B_{k}\rangle.$
Since $(B(E(t)))_{t\geq 0}$ is a martingale, we have $\mathbb{E}_{B}[\Delta B_{k}]=0$. Taking $\mathbb{E}_{B}$ on both sides and noting that $\mathbb{E}_{B}[|\Delta B_{k}|^{2}]=\Delta E_{k}$, we obtain
$\mathbb{E}_{B}[|X_{k+1}|^{2}]$ $=\mathbb{E}_{B}[|X_{k}|^{2}]+\mathbb{E}_{B}[|f(X_{k})|^{2}]|\Delta E_{k}|^{2}+\mathbb{E}_{B}[|g(X_{k})|^{2}]\Delta E_{k}$
$\quad+2\mathbb{E}_{B}[\langle X_{k},f(X_{k})\rangle]\Delta E_{k}.$
Applying Assumptions 8 and 9, we arrive at the fundamental one-step inequality
$\mathbb{E}_{B}[|X_{k+1}|^{2}]$ $\leq(1-L\Delta E_{k}+K^{2}(\Delta E_{k})^{2})\mathbb{E}_{B}[|X_{k}|^{2}].$ (12)
The challenge in the subsequent analysis is that directly adding $(\Delta E_{k})^{2}$ into the exponent would yield a positive exponential term $\sum(\Delta E_{k})^{2}$, which could potentially dominate the polynomial decay induced by $L\sum\Delta E_{k}$. To circumvent this difficulty, we employ a localization argument based on stopping times.
For each $m\in\mathbb{N}$, define the stopping time
$\tau_{m}:=\inf\{t\geq 0:E(t)\geq m\}\quad(m\in\mathbb{N}),$
which increases to $\infty$ as $m\to\infty$ and satisfies $\tau_{m}<\infty$ almost surely. Since $E(\cdot)$ is continuous and bounded on the compact interval $0,\tau_{m}]$, its modulus of continuity
$\omega_{m}(\Delta t):=\sup_{0\leq s\leq\tau_{m},\,0\leq h\leq\Delta t}\big(E(s+h)-E(s)\big)$
satisfies $\omega_{m}(\Delta t)\to 0$ as $\Delta t\to 0$ by uniform continuity on compact sets. Consequently, for all $k$ such that the event $\{t_{k+1}\leq\tau_{m}\}$ holds, we have
$(\Delta E_{k})^{2}\leq\omega_{m}(\Delta t)\Delta E_{k}.$ (13)
Substituting (13) into (12) and using the elementary inequality $1+x\leq e^{x}$, we obtain
$\mathbb{E}_{B}\left[|X_{k+1}|^{2}\mathbf{1}_{\{t_{k+1}\leq\tau_{m}\}}\right]$ $\leq\exp\big(-(L-K^{2}\omega_{m}(\Delta t))\Delta E_{k}\big)\mathbb{E}_{B}\left[|X_{k}|^{2}\mathbf{1}_{\{t_{k}\leq\tau_{m}\}}\right].$
Iterating this inequality from $k=0,\ldots,n-1$ yields
$\mathbb{E}_{B}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]\leq|X_{0}|^{2}\exp\big(-(L-K^{2}\omega_{m}(\Delta t))E(t_{n}\wedge\tau_{m})\big).$ (14)
On the event $\{t_{n}\leq\tau_{m}\}$, we have $t_{n}\wedge\tau_{m}=t_{n}$. Taking $\mathbb{E}_{D}$ on both sides of (14) and using $\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\leq 1$, we obtain
$\mathbb{E}\!\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]\leq|X_{0}|^{2}\,\mathbb{E}_{D}\!\left[\exp\!\big(-(L-K^{2}\omega_{m}(\Delta t))\,E(t_{n})\big)\right].$
Since $\omega_{m}(\Delta t)\to 0$ as $\Delta t\to 0$ for each fixed $m$, there exists $N(m)\in\mathbb{N}$ such that for all $n\geq N(m)$,
$K^{2}\omega_{m}(\Delta t)\leq\frac{L}{2}.$ (15)
Then, by applying (15), we obtain for all $n\geq N(m)$,
$\mathbb{E}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]$ $\leq\mathbb{E}\left[|X_{0}|^{2}\right]\mathbb{E}_{D}\left[e^{-\frac{L}{2}E(t_{n})}\right].$ (16)
By Lemma 11, taking $\lambda=\frac{L}{2}$ and letting $n\to\infty$, we deduce that for any fixed $m$,
$\limsup_{n\to\infty}\frac{\log\mathbb{E}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right]}{\log t_{n}}$ $\leq\limsup_{n\to\infty}\frac{\log\mathbb{E}_{D}\left[e^{-\frac{L}{2}E(t_{n})}\right]}{\log t_{n}}=-\alpha.$
To remove the localization, we invoke the monotone convergence theorem. Since $\tau_{m}\to\infty$ almost surely and $\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\to 1$ as $m\to\infty$, we have
$\mathbb{E}[|X_{n}|^{2}]=\lim_{m\to\infty}\mathbb{E}\left[|X_{n}|^{2}\mathbf{1}_{\{t_{n}\leq\tau_{m}\}}\right].$
Consequently,
$\limsup_{n\to\infty}\frac{\log\mathbb{E}[|X_{n}|^{2}]}{\log t_{n}}\leq-\alpha,$
which completes the proof. ∎
5 Numerical simulations
In this section, we present numerical experiments to validate our theoretical results. We first define our error metric and then apply our numerical scheme to several representative examples.
To quantify the accuracy of our numerical scheme, we employ the $L_{1}$ error metric defined as
$e_{i}=\mathbb{E}\left|X(T)-X_{\Delta t_{i}}(T)\right|,$
where $X_{\Delta t_{i}}(T)$ represents the numerical solution at terminal time $T$ computed with step size $\Delta t_{i}=2^{-i}$, and $X(T)$ denotes the true solution. For our experiments, we set $T=1$ and approximate the expectation via Monte Carlo simulation with $M=10^{4}$ independent realizations
$e_{i}$ $\approx\frac{1}{M}\sum_{j=1}^{M}\left|X^{j}(T)-X^{j}_{\Delta t_{i}}(T)\right|.$
We employ a sufficiently fine discretization with step size $\Delta t=2^{-15}$ as the reference solution. The convergence analysis is conducted by computing numerical solutions with a sequence of decreasing step sizes $(2^{3}\Delta t,2^{4}\Delta t,2^{5}\Delta t,2^{6}\Delta t)$. The corresponding convergence rates are then determined through linear regression in logarithmic coordinates, specifically by fitting $\log(e_{i})$ against $\log(\Delta t_{i})$.
For the stability analysis, we simulate $10^{4}$ sample paths up to the terminal time $T=40$ with a step size of $\Delta t=2^{-13}$ and stability index $\alpha=0.6,0.7,0.8$. This fine discretization allows us to accurately capture the long-term behavior of the example while ensuring numerical stability according to the conditions established in Theorems 10 and 12.
Example 1.
We consider a two-dimensional coupled time-changed SDE
$\left\{\begin{aligned} dX_{1}(t)&=-(X_{1}(t)+X_{2}(t))\,dE(t)+2(X_{1}(t)+X_{2}(t))\,dB_{1}(E(t)),\\ dX_{2}(t)&=-2(X_{1}(t)+X_{2}(t))\,dE(t)+(X_{1}(t)+X_{2}(t))\,dB_{2}(E(t)),\end{aligned}\right$.
with initial conditions $X_{1}(0)=1$ and $X_{2}(0)=2$. Here, $B_{1}$ and $B_{2}$ are independent standard Brownian motions.
This example presents several interesting features.
  • The dynamics are coupled, with each component affecting the evolution of the other
  • The example is driven by two independent Brownian motions
  • Both drift and diffusion terms have multiplicative noise structure
One can easily verify that this example satisfies all conditions of Theorem 7. We calculate the numerical convergence rates using the method described above. Table 1 displays the empirical convergence rates for different values of the parameter $\alpha$. The numerical results for this coupled example align well with our theoretical predictions, thus providing strong support for the validity of our analytical framework. Additionally, we provide a comparison of the convergence rates for $\alpha=0.6,0.8$ in Figure 2, which visually demonstrates the agreement between theoretical and numerical results.
Figure 2: The left figure shows the convergence rates for $\alpha=0.6$, and the right figure shows the convergence rates for $\alpha=0.8$. The blue line represents the theoretical convergence rates, and the red line represents the numerical convergence rates.
To verify the observation in the remark following Theorem 7, we add a drift coefficient of $1$ to the subordinator. Figure 3 shows the numerical convergence rates for $\alpha=0.6$ and $\alpha=0.8$ in this setting, which approach the classical rate of $1/2$.
Figure 3: Convergence rates with positive linear drift coefficient 11 for $\alpha=0.6$ (left) and $\alpha=0.8$ (right).
Table 1: Convergence rates with different $\alpha$ values.
$\alpha$ 0.6000 0.6500 0.7000 0.7500 0.8000 0.8500 0.9000
Theoretical rate 0.4000 0.4125 0.4250 0.4375 0.4500 0.4625 0.4750
Numerical rate 0.4005 0.4069 0.4200 0.4349 0.4474 0.4737 0.4870
Example 2.
We consider a scalar time-changed SDE
$dX(t)=-\frac{1}{3}X(t)\,dE(t)+\frac{1}{3}X(t)\,dB(E(t)),$
where $B$ is a standard Brownian motion.
For this example, we can easily verify that the Assumptions 8 and 9 are satisfied. According to Theorem 10, the true solution of this time-changed SDE exhibits asymptotic polynomial stability in the mean square sense, with the decay rate $\alpha$. As demonstrated in [4], the explicit analytical solution of this time-changed SDE can be expressed in the following closed form
$X(t)=X(0)e^{-\frac{7}{18}E(t)+\frac{1}{3}B(E(t))}.$
Figures 5 and 5 present the mean square trajectories of the true solution and the EM method solution for different values of the stability index $\alpha\in\{0.6,0.7,0.8\}$. The influence of the stability index on the decay rate is apparent, with smaller values of $\alpha$ resulting in slower decay, consistent with our theoretical analysis.
Figure 4: Mean square trajectories of the true solution for different values of $\alpha$.
Figure 5: Mean square trajectories of the EM method solution for different values of $\alpha$.
To rigorously verify the asymptotic decay rate quantitatively, we employ a normalization procedure as illustrated in Figures 7 and 7. The reference curves $t^{-\alpha}$ are normalized at $t=32$ to align with the corresponding mean square trajectories of both the true solution and the EM method solution at this reference point. The comparative analysis over the interval $t\in[32,40]$ reveals a high degree of concordance between each pair of curves. This precise alignment between theoretical predictions and numerical simulations across different values of $\alpha$ strongly validates the asymptotic stability properties established in Theorems 10 and 12 and confirms the role of the stability index in determining the polynomial decay rate. Furthermore, it demonstrates the capacity of our equidistant-step EM method to accurately preserve the stability characteristics of the time-changed SDEs, a critical property for numerical methods applied to non-Markovian systems with memory effects.
Figure 6: Comparison between the mean square trajectory of the true solution and the reference curve $t^{-\alpha}$ over the interval $t\in[32,40]$.
Figure 7: Comparison between the mean square trajectory of the EM method solution and the reference curve $t^{-\alpha}$ over the interval $t\in[32,40]$.
These two examples demonstrate the strong convergence property and the stability, respectively, of our proposed equidistant-step EM method for time-changed SDEs, providing comprehensive numerical validation of the theoretical results established in this paper.
Appendix A Proof of Lemma 3
Proof.
We begin by expressing the $n$-th power of the integral as a multiple integral. Applying the change of variable formula for multiple integrals, we obtain
$\left(\int_{a}^{b}\mathrm{d}E(t)\right)^{n}=n\int_{a}^{b}\left(\int_{a}^{t_{n}}\mathrm{d}E(t_{n-1})\right)^{n-1}\mathrm{d}E(t_{n}).$
Iterating this procedure $n$ times yields
$\left(\int_{a}^{b}\mathrm{d}E(t)\right)^{n}=n!\int_{a}^{b}\int_{a}^{t_{n}}\cdots\int_{a}^{t_{2}}\mathrm{d}E(t_{1})\mathrm{d}E(t_{2})\cdots\mathrm{d}E(t_{n}),$
where the integration domain satisfies the ordering constraint $a<t_{1}<t_{2}<\cdots<t_{n}<b$.
To evaluate the expectation of this multiple integral, we invoke the theory of Cox processes. Let us define the random measure $\Pi$ on $0,\infty)$ by $\Pi((s,t])=E(t)-E(s)$ for $t>s\geq 0$. Consider a Cox process $\{C(t)\}_{t\geq 0}$ driven by $\Pi$, that is, conditional on $\Pi=\lambda$, the process $\{C(t)\}$ follows the distribution of an inhomogeneous Poisson process with intensity $\lambda$. As established in [26], $\{C(t)\}$ is a renewal process with renewal function
$u(t)=\mathbb{E}[C(t)]=\mathbb{E}[E(t)]=\frac{t^{\alpha}}{\Gamma(\alpha+1)}.$
For a renewal process $C(t)$, following [27], we have
$\mathbb{E}[\mathrm{d}C(t_{1})\mathrm{d}C(t_{2})\cdots\mathrm{d}C(t_{n})]=\prod_{i=1}^{n}u^{\prime}(t_{i}-t_{i-1})\mathrm{d}t_{i},$
where $t_{0}=0$. Since the factorial moments of the Cox process $C(t)$ equal the ordinary moments of its driving measure $\Pi$ (see [27]), we obtain
$\mathbb{E}[\mathrm{d}E(t_{1})\mathrm{d}E(t_{2})\cdots\mathrm{d}E(t_{n})]=\prod_{i=1}^{n}u^{\prime}(t_{i}-t_{i-1})\mathrm{d}t_{i}.$
Thus, denoting $I=\mathbb{E}\left[\left(\int_{a}^{b}\mathrm{d}E(t)\right)^{n}\right]$, we have
$I$ $=n!\int_{a}^{b}\int_{a}^{t_{n}}\int_{a}^{t_{n-1}}\cdots\int_{a}^{t_{2}}\mathbb{E}\left[\mathrm{d}E(t_{1})\cdots\mathrm{d}E(t_{n-1})\mathrm{d}E(t_{n})\right]$
$=\frac{n!\alpha^{n}}{\Gamma^{n}(\alpha+1)}\int_{a}^{b}\int_{a}^{t_{n}}\int_{a}^{t_{n-1}}\cdots\int_{a}^{t_{2}}\prod_{j=1}^{n}(t_{j}-t_{j-1})^{\alpha-1}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n-1}\mathrm{d}t_{n},$
where $t_{0}=0$ for the general integral formula.
We now analyze the nested integrals systematically. To facilitate the analysis, let us introduce the change of variables $t_{j}=a+s_{j}(b-a)$ for $j=1,2,\ldots,n$, which maps the interval $[a,b]$ to $[0,1]$. Under this transformation, we have $dt_{j}=(b-a)ds_{j}$ and the integration domain becomes $0<s_{1}<s_{2}<\cdots<s_{n}<1$. The innermost integral transforms to
$I_{1}$ $=\int_{a}^{t_{2}}(t_{2}-t_{1})^{\alpha-1}t_{1}^{\alpha-1}\mathrm{d}t_{1}$
$=\int_{0}^{s_{2}}((a+s_{2}(b-a))-(a+s_{1}(b-a)))^{\alpha-1}(a+s_{1}(b-a))^{\alpha-1}(b-a)\mathrm{d}s_{1}$
$=(b-a)^{\alpha}\int_{0}^{s_{2}}(s_{2}-s_{1})^{\alpha-1}(a+s_{1}(b-a))^{\alpha-1}\mathrm{d}s_{1}.$
Since $a>0$ and $\alpha<1$, the function $(a+s_{1}(b-a))^{\alpha-1}$ is monotonically decreasing with respect to $s_{1}$ on $[0,1]$. Therefore,
$I_{1}$ $\leq(b-a)^{\alpha}\int_{0}^{s_{2}}(s_{2}-s_{1})^{\alpha-1}a^{\alpha-1}\mathrm{d}s_{1}$
$=a^{\alpha-1}(b-a)^{\alpha}\int_{0}^{s_{2}}(s_{2}-s_{1})^{\alpha-1}\mathrm{d}s_{1}.$
Making the substitution $w=s_{2}-s_{1}$, we get
$I_{1}$ $\leq a^{\alpha-1}(b-a)^{\alpha}\int_{0}^{s_{2}}w^{\alpha-1}\mathrm{d}w$
$=\frac{a^{\alpha-1}(b-a)^{\alpha}s_{2}^{\alpha}}{\alpha}$
$\leq\frac{a^{\alpha-1}(b-a)^{\alpha}}{\alpha},$
since $s_{2}\leq 1$.
Consequently, we have
$I$ $\leq C_{1}(b-a)^{\alpha}\int_{a}^{b}\int_{a}^{t_{n}}\int_{a}^{t_{n-1}}\cdots\int_{a}^{t_{3}}\prod_{j=3}^{n}(t_{j}-t_{j-1})^{\alpha-1}\mathrm{d}t_{2}\cdots\mathrm{d}t_{n-1}\mathrm{d}t_{n},$
where $C_{1}>0$ is a constant depending on $a$ and $\alpha$ but independent of bab-a.
Proceeding similarly with the substitution $t_{j}=a+s_{j}(b-a)$ for each nested integral, for $k=2,3,\ldots,n-1$, we obtain
$\int_{a}^{t_{k+1}}(t_{k+1}-t_{k})^{\alpha-1}\mathrm{d}t_{k}\leq C_{k}(b-a)^{\alpha},$
where each $C_{k}>0$ depends on $a$ and $\alpha$ but is independent of bab-a.
By iterating this process n1n-1 times and applying the change of variables to all integrals, we obtain
$I$ $\leq C(b-a)^{(n-1)\alpha}\int_{a}^{b}1\mathrm{d}t_{n}$
$=C(b-a)^{1+(n-1)\alpha},$
where $C=C_{1}C_{2}\cdots C_{n-1}$ is a positive constant depending on $a$ and $\alpha$ but independent of bab-a. This completes the proof. ∎
Appendix B Proof of Theorem 10
Proof.
Let us apply the time-changed Itô formula to $e^{LE(s)}|X(s)|^{2}$ where $L>0$ is the constant from Assumption 9.
$d(e^{LE(s)}|X(s)|^{2})$ $=Le^{LE(s)}|X(s)|^{2}\,dE(s)+e^{LE(s)}\,d|X(s)|^{2}$
$=Le^{LE(s)}|X(s)|^{2}\,dE(s)+2e^{LE(s)}\langle X(s),f(X(s))\rangle\,dE(s)$
$\quad+e^{LE(s)}|g(X(s))|^{2}\,dE(s)+2e^{LE(s)}\langle X(s),g(X(s))\rangle\,dB(E(s))$
$=e^{LE(s)}\left(L|X(s)|^{2}+2\langle X(s),f(X(s))\rangle+|g(X(s))|^{2}\right)\,dE(s)$
$\quad+e^{LE(s)}2\langle X(s),g(X(s))\rangle\,dB(E(s)).$
By Assumption 9, we have $2\langle X(s),f(X(s))\rangle+|g(X(s))|^{2}\leq-L|X(s)|^{2}$, which gives us
$d(e^{LE(s)}|X(s)|^{2})$ $\leq e^{LE(s)}\left(L|X(s)|^{2}-L|X(s)|^{2}\right)\,dE(s)$
$\quad+e^{LE(s)}2\langle X(s),g(X(s))\rangle\,dB(E(s))$
$=e^{LE(s)}2\langle X(s),g(X(s))\rangle\,dB(E(s)).$
Since $\mathbb{E}_{B}\left[\int_{0}^{t}e^{LE(s)}2\langle X(s),g(X(s))\rangle\,dB(E(s))\right]=0$, taking the conditional expectation $\mathbb{E}_{B}[\cdot]$ on both sides and integrating from 0 to $t$, we get
$\mathbb{E}_{B}[e^{LE(t)}|X(t)|^{2}]\leq|X(0)|^{2}.$
Rearranging, we obtain
$\mathbb{E}_{B}[|X(t)|^{2}]\leq|X(0)|^{2}\cdot e^{-LE(t)}.$
Taking the expectation $\mathbb{E}_{D}$ on both sides, we derive
$\mathbb{E}[|X(t)|^{2}]\leq|X(0)|^{2}\cdot\mathbb{E}_{D}\left[e^{-LE(t)}\right].$
Taking the logarithm of both sides and dividing by $\log t$, we have
$\frac{\log\mathbb{E}[|X(t)|^{2}]}{\log t}\leq\frac{\log|X(0)|^{2}}{\log t}+\frac{\log\mathbb{E}_{D}\left[e^{-LE(t)}\right]}{\log t}.$
As $t\to\infty$, the first term on the right-hand side approaches zero. For the second term, by Lemma 11 and noting that $L>0$ under Assumption 9, we have
$\limsup_{t\to\infty}\frac{\log\mathbb{E}_{D}\left[e^{-LE(t)}\right]}{\log t}=-\alpha.$
Therefore, we conclude that
$\limsup_{t\to\infty}\frac{\log\mathbb{E}[|X(t)|^{2}]}{\log t}\leq-\alpha,$
which establishes the asymptotic polynomial stability of the true solution with the precise decay rate of $\alpha$. ∎